Glucose has the formula C6H12O6. What is the maximum mass of carbon that could be obtained from 200.00g of glucose?
1
Expert's answer
2020-05-05T03:40:08-0400

The molecular weight of glucose equals:

Mr(C6H12O6) = (12 × 6) + (12 × 1) + (16 × 6) = 180 g/mol

As a glucose molecule contains six carbons, the molecular weight of those carbons equals:

Mr(C6) = 72 g/mol

As 180 g/mol corresponds to 200.00 g of glucose and 72 g/mol corresponds to the unknown mass of carbon in glucose:

180 g/mol / 72 g/mol = 200.00 g / m(carbon)

From here:

m(carbon) = 72 g/mol × 200.00 g / 180 g/mol = 80 g


Answer: 80 g